# Unveiling Attacks Ahead of Exploit: adversarial Contract Detection Based on DeFi Attack Features

### 14 Jan 2024 v1

@@ 4.1.3 Verification Status
Contracts can be verified by providing their source code and compiler settings to platforms like Etherscan. Verified contracts are open source, allowing anyone to view their source code. Many development frameworks, such as Foundry[22], offer options to verify contracts upon deployment. Generally, benign contracts tend to be verified to demonstrate their security and trustworthiness. Attackers, on the other hand, typically do not verify the adversarial contracts they deployed, exposing their malicious intent.
4.1.4 Implemented Function
@@
This category covers features related to functions implemented in the contract. It includes the number of public functions, private functions, total functions, and the proportion of public functions. In particular, we incorporate the number of flashloan callback functions and their proportion among all public functions. This choice is motivated by the fact that many attack scenarios may require significant token holdings, and attackers may not possess these tokens, necessitating the use of flashloans to borrow tokens and execute complex attack strategies within corresponding callback functions. Additionally, since lending pool sizes may be limited, attackers might need to borrow tokens from multiple sources, resulting in multiple flashloan callback functions. The flashloan callback functions are referred to in Table 1.
4.1.5 Function Call
@@
The Function Call provides insights into the specific calls of each function in the contract, including the number of external calls and internal calls, the total number of calls, and the ratio of external calls. It’s important to clarify that while we refer to them as internal calls, they are distinct from external calls, as they utilize the jump opcode instead of the call opcode. We also include the number of delegatecall.
@@
In addition, we consider external calls related to ERC-20 tokens, including their count and proportion. The specific functions are outlined in Table 1. Generally, the objective of DeFi attack is to gain economic benefits, i.e., acquire valuable tokens and eventually convert them into native tokens for withdrawal. Therefore, external calls within adversarial contracts often involve token-related functions.
@@
While the withdraw and deposit functions are not ERC-20 standard functions, they are implemented for wrapped native tokens. Among the functions within DEX implementations, In addition to swap related functions, we also include some UniswapV2 protocol functions, such as sync and skim, as they are also common functions used in some attacks.
@@
Furthermore, We introduce two additional features, MaxTokenCall and AvgTokenCall. MaxTokenCall represents the maximum number of token-related calls within a single public function execution flow, which helps identify functions with exceptionally high token interactions. AvgTokenCall, on the other hand, indicates the average number of token-related calls per public function execution flow, providing a measure of the average token interaction level.
4.1.6 Contract Type
Token contracts and proxy contracts are widely prevalent on the blockchain, while adversarial contracts typically do not fall into these categories. Additionally, some attackers have a tendency to self-destruct their contracts subsequent to the completion of the attack. As a result, we include the contract type as one of the distinctive features.
4.2 Feature Extraction
4.2.1 Transaction Data Retrieval
@@
4.2.2 Explorer Data Retrieval
@@
We employ Algorithm 1 to obtain the labels of the fund source. We begin by retrieving the complete transaction history of the contract creator, encompassing both normal and internal transactions. We then identify the earliest transaction where native tokens were transferred to the contract creator’s address, considering its sender as the contract creator’s fund source.
The obtained address at this stage cannot serve as a feature directly. Therefore, we combine the address with the publicly available address labels and tags from Etherscan and Bscscan. For each fund source address corresponding to our contract samples, we construct a mapping table associating addresses with their respective labels. This mapping table enables us to extract explicit fund source features for all fund source addresses.
We categorize fund source addresses into five types:
@@
    Unknown: Addresses lacking specific labels but possessing a substantial history of transfer transactions. We cannot categorize them accurately.
  5. 5)
    Transfer: Addresses with minimal transaction activity, suggesting that the fund source is an ordinary address used for basic transfer.
Input: The contract creator $c$
Output: The fund source label of $c$
1 init
2       $C_{\textnormal{hop}}\leftarrow$ 2
3       $C_{\textnormal{threshold}}\leftarrow$ 10000
4       $D_{\textnormal{label}}\leftarrow$label1: [address1, address2], label2: [address3, address4], …}
5
6
7while $C_{\textnormal{hop}}$geq 0 do
8       $T_{\textnormal{xn}}\leftarrow$ First normal transaction of $c$ with a native token transfer in.
9       $T_{\textnormal{xi}}\leftarrow$ First internal transaction of $c$ with a native token transfer in.
10       $T_{\textnormal{x}}\leftarrow$ The transaction with the smaller blocknum between $T_{\textnormal{xn}}$ and $T_{\textnormal{xi}}$.
11       $f_{\textnormal{address}}\leftarrow$ The sender address of $T_{\textnormal{x}}$.
12       if $f_{\textnormal{address}}$ in $D_{\textnormal{label}}$ then
13             return The label of $f_{\textnormal{address}}$ in $D_{\textnormal{label}}$
14       end if
15      $C_{\textnormal{hop}}$ = $C_{\textnormal{hop}}$ 1;
16       if $C_{\textnormal{hop}}$ ==absent$=$ = 0 then
17             $n\leftarrow$ Historical transaction count of $f_{\textnormal{address}}$
18             if $n$ < $C_{\textnormal{threshold}}$ then
19                   return’Transfer’
20             end if
21            return’Unknown’
22       end if
23      $c\leftarrow f_{\textnormal{address}}$
24 end while
Algorithm 1 Fund Source Labeling Algorithm
To further refine the accuracy of fund source feature labels, in the case where we identify the fund source address as belonging to the Transfer category, we will continue to trace it once, that is, conduct an additional lookup to ascertain the label of the fund source corresponding to the fund source address again. This step is crucial to prevent attackers from using intermediary address transfers to evade detection of their malicious fund source, thereby ensuring more precise identification of the fund source.
4.2.3 Decompilation
@@
Transaction features Logistic Regression 0.9804 0.5053 0.9057 0.6486 0.0181
Decision Trees 0.9845 0.6667 0.4528 0.5393 0.0046
Random Forest 0.9883 0.7500 0.6226 0.6804 0.0042
MLP 0.9804 0.5059 0.8113 0.6232 0.0162
LightGBM 0.9872 0.7111 0.6038 0.6531 0.0050
Contract features Logistic Regression 0.9453 0.2629 0.9623 0.4130 0.0550
Decision Trees 0.9879 0.6721 0.7736 0.7193 0.0077
Random Forest 0.9917 0.7925 0.7925 0.7925 0.0042
MLP 0.9876 0.6429 0.8491 0.7317 0.0096
LightGBM 0.9925 0.8367 0.7736 0.8039 0.0031
Hybrid features Logistic Regression 0.9872 0.6234 0.9057 0.7385 0.0112
Decision Trees 0.9925 0.8511 0.7547 0.8000 0.0027
Random Forest 0.9970 0.9245 0.9245 0.9245 0.0015
MLP 0.9951 0.8448 0.9245 0.8829 0.0035
LightGBM 0.9977 0.9434 0.9434 0.9434 0.0012
@@
Table 3: Performance Metrics for Different Features and Models
Finally, we utilize tools provided by Dedaub[17] to decompile the contract bytecode, allowing us to analyze and extract the remaining features. Regardless of whether the contract is verified, to facilitate subsequent unified processing, we will perform the decompilation.
Specifically, we first decompile the contract’s bytecode (transaction input data) and runtime code (deployed on the chain) to obtain pseudocode. Then, we use regular expressions to extract information about implemented functions, including their names and visibility, thereby obtaining all features for the Implemented Function category. Subsequently, we also utilize regular expressions to further analyze the function calls within each function, obtaining features for the Function Call category, and determine whether the contract can self-destruct.
It’s worth noting that Dedaub’s decompilation attempts to automatically resolve function signatures, but the results may not always be complete. Therefore, we use both function names and function selectors to detect the implementation of flashloan callback functions and token-related calls.
To extract the MaxTokenCall and AvgTokenCall features, we analyze the execution flow of each public function, which means detecting the internal calls and potential flashloan function calls, and then counting token related calls involved in them.
Furthermore, to determine whether the contract is a token contract, we inspect the implemented function to check if it fully implements all functions of a specific token ERC standard. Specifically, we implement detection for ERC-20[1], ERC-721[5], and ERC-1155[4] token contract standards. (Other token ERC standards ERC-777[3] and ERC-4626[9] are backward-compatible with ERC-20)
@@
For identifying proxy contracts, despite various proxy patterns, most proxy contracts, following ERC-1967[6], typically saves the logic contract’s address in a specific storage slot. Therefore, if a contract involves a delegatecall and the contract creation code contains the slot value, we consider it a proxy contract, which covers the majority of cases.
4.3 Classifier Building
@@
For the fund source feature, which consists of categorical data with low cardinality, we applied one-hot encoding. Additionally, to ensure that features of different dimensions are more comparable in numerical terms and to enhance the classifier’s accuracy, we standardized the data using Z-score normalization. This process centers each feature’s values around 0 with a standard deviation of 1.
@@
Finally, when selecting an appropriate supervised machine learning algorithm to construct the classifier, we explore various methods. These methods include classical approaches such as Logistic Regression and Decision Trees. We also try ensemble learning approaches, such as Random Forest based on bagging and the LightGBM algorithm based on boosting. In addition, we considered the use of a classic neural network architecture, the Multilayer Perceptron (MLP), as one of the candidate models.
5 Evaluation
5.1 Metrics
@@
In addition to commonly used binary classification performance metrics such as accuracy, precision, recall, F1-Score, we also calculate the False Positive Rate (FPR) metric, which holds particular significance in our scenario. A high FPR implies that the classifier incorrectly identifies benign contracts as adversarial, leading to substantial manual audit costs.
@@
A distinguishing aspect of our research compared to traditional supervised machine learning problems is the explicit chronological relationship between contract deployment. Consequently, traditional cross-validation techniques may not be suitable for our dataset, as they assume that each data is independent. And our objective is to train a model based on existing data to detect potential future attacks, so for sustainability reasons, we order the contract samples chronologically, allocating the first 80% for training and the remaining portion for testing to calculate relevant metrics, instead of relying on cross-validation.
5.2 Performance
@@
Table 3 presents our test results. Overall, the LightGBM model outperforms others, demonstrating the best performance across all metrics when using all features. It exhibits a remarkably high detection rate for adversarial samples, with an F1-Score of 0.9434, while maintaining an exceptionally low FPR of 0.12%. Other models also perform well, indicating the applicability of our approach to the dataset.
To the best of our knowledge, this is the first academic work on the detection of adversarial contracts utilized in DeFi attack. The only similar research made public so far is by the Forta Network[21], which also uses machine learning. However, their approach differs as it focuses solely on low-level bytecode. We used their method as a baseline for comparison.
Model Precision Recall F1-Score
forta 0.8778 0.5520 0.6208
ours 0.9434 0.9434 0.9434
Table 4: Model Performance Metrics
As shown in Table 4, our model outperforms Forta’s most recently published model (released on February 6, 2023) across all performance metrics. This highlights the superiority of our method.
5.3 Feature Importance

### 02 Feb 2024 v2

@@ 4.1.3. Verification Status
Contracts can be verified by providing their source code and compiler settings to platforms like Etherscan. Verified contracts are open source, allowing anyone to view their source code. Many development frameworks, such as Foundry (foundry, 36), offer options to verify contracts upon deployment. Generally, benign contracts tend to be verified to demonstrate their security and trustworthiness. Attackers, on the other hand, typically do not verify the adversarial contracts they deployed, exposing their malicious intent. This can also be corroborated by our contract samples presented in Table 3.
4.1.4. Implemented Function
Figure 6. Distribution of public function number and private function number for both adversarial and benign contract samples.
This category covers features related to implemented functions in the contract, such as the number of public functions, private functions, total functions, and the proportion of public functions. Upon analyzing our contract samples, as illustrated in Figure 6, it becomes apparent that benign contracts frequently incorporate more than 20 public functions, while adversarial contracts tend to implement fewer public functions. Furthermore, adversarial contracts tend to implement fewer private functions.
Figure 7. Distribution of implemented flashloan callback function number for both adversarial and benign contract samples.
@@ 4.1.4. Implemented Function
In particular, we incorporate the number of flashloan callback functions and their proportion among all public functions. The flashloan callback functions are referred to in Table 4. This choice is motivated by the fact that many attack scenarios may require significant token holdings, and attackers may not possess these tokens, necessitating the use of flashloans to borrow tokens and execute complex attack strategies within corresponding callback functions. Additionally, since lending pool sizes may be limited, attackers might need to borrow tokens from multiple sources, resulting in multiple flashloan callback functions. In our contract samples, as shown in Figure 7, it is noticeable that approximately two-thirds of adversarial contracts implement flashloan callback functions, while almost all benign contracts do not involve such callbacks.
Flashloan Callback ERC-20 DEX
uniswapV2Call
totalSupply
uniswapV3FlashCallback
balanceOf
uniswapV3SwapCallback
transfer
swap
pancakeCall
allowance
sync
pancakeV3FlashCallback
approve
skim
DPPFlashLoanCall
transferFrom
getReserves
DVMFlashLoanCall
mint
token0
executeOperation
burn
token1
receiveFlashLoan
withdraw
onFlashLoan
deposit
BiswapCall
Table 4. Flashloan Callback and Token-related Function
4.1.5. Function Call
@@
This category provides insights into the specific calls of each function in the contract, including the number of external and internal calls, the total number of calls, the ratio of external calls, and the number of delegatecall. It’s important to clarify that while we refer to them as internal calls, they are distinct from external calls because they utilize the jump opcode instead of the call opcode, signifying a fundamental difference in the underlying EVM execution. In our contract samples, as shown in Figure 8, over half of adversarial contracts involve more than 20 external calls, while most benign contracts have fewer than 10 external calls. Additionally, benign contracts tend to have more internal calls compared to adversarial contracts.
@@
We have also taken into consideration external calls related to ERC-20 tokens, including their count and proportion. The specific functions are outlined in Table 4. Generally, the objective of DeFi attack is to gain financial benefits, i.e., acquire valuable tokens and eventually convert them into native tokens for withdrawal. Therefore, external calls within adversarial contracts often involve token-related functions.
@@
While the withdraw and deposit functions are not ERC-20 standard functions, they are implemented for wrapped native tokens. Among the functions within DEX implementations, in addition to swap related functions, we also include some UniswapV2 (UniswapV2, 63) protocol functions, such as sync and skim, as they are also common functions used in some attacks. In our sample contracts, as shown in Figure 9, it is evident that over 70% of benign contracts do not have token-related calls. In contrast, almost all adversarial contracts involve token-related calls, with over half of them having 16 or more token-related calls.
@@
Furthermore, we introduce two novel features to assess function interactions with tokens: MaxTokenCall and AvgTokenCall. To better understand their significance, it’s essential to first consider the concept of the function execution flow. The flow encompasses the complete process a function undergoes during its operation. Specifically, when a public function is executed, it often involves invoking other implemented functions. MaxTokenCall is designed to capture the peak level of token-related interactions within a single public function execution flow, which helps identify functions with exceptionally high token interactions. On the other hand, AvgTokenCall indicates the average number of token-related calls per public function execution flow, providing a measure of the average token interaction level.
Figure 8. Distribution of external and internal call number for both adversarial and benign contract samples.
Figure 9. Distribution of token-related call number for both adversarial and benign contract samples.
4.1.6. Contract Type
Type Sample YES NO
Token benign 9245 (71.1%) 3755 (28.9%)
adversarial 0 (0%) 269 (100%)
Proxy benign 735 (5.7%) 12265 (94.3%)
adversarial 0 (0%) 269 (100%)
Selfdestruct benign 78 (0.6%) 12922 (99.4%)
adversarial 36 (13.4%) 233 (86.6%)
Table 5. Distribution of benign and adversarial contracts across different contract types.
As previously mentioned in §2.2, tokens are fundamental to DeFi. Consequently, token contracts are extensively prevalent on the blockchain, representing one of the primary types of contracts. Similarly, proxy contracts are also widely used on the blockchain, serving a different but equally vital role. They act as intermediaries to forward function calls to other contracts, allowing for flexible updates to the contract’s logic without changing the proxy contract’s address, which is important for upgradability and flexibility in smart contracts. While adversarial contracts typically do not fall into either of these types. Furthermore, the EVM provides the selfdestruct instruction, enabling a contract to delete itself from the blockchain. Some attackers have a tendency to self-destruct their contracts subsequent to the completion of the attack. As a result, we include the contract type as distinctive features.
According to our contract type identification algorithm (detailed in §4.2), in our contract samples, as shown in Table 5, none of the adversarial contracts fall into the token or proxy contract types. In contrast, over 70% of benign samples are token contracts, and proxy contracts make up more than 5%. Additionally, less than 1% of benign contracts use the selfdestruct instruction, while over 10% of adversarial contracts do so.
4.2. Feature Extraction
4.2.1. Transaction Data Retrieval
@@
4.2.2. Explorer Data Retrieval
@@
[…]
Algorithm 1 Fund Source Labeling Algorithm
[…]
13:     end if
[…]
15:     if $n\geq 10000$ then
16:         return ’Unknown’
17:     end if
[…]
20:         return ’Unknown’
21:     end if
[…]
23:end while
[…]
Specifically, We categorize fund source addresses into four types:
@@
    Unknown: Addresses lacking specific labels but possessing a substantial historical transaction volume. We cannot categorize them accurately.
Figure 10. Distribution of fund source labels for both adversarial and benign contract samples.
Based on the fund source labeling algorithm, we conducted a statistical analysis for our contract samples. As depicted in Figure 10, in the benign samples, more than two-thirds of the fund source labels are identified as Safe, approximately one-quarter are labeled as Unknown, and a small portion fell into the categories of Anonymous and Bridge. Conversely, in the adversarial samples, the majority of labels are categorized as Anonymous, with a small proportion belong to the Safe category.
4.2.3. Decompilation
Finally, we employ Dedaub’s tool (Dedaub, 23) to decompile the contract bytecode, allowing us to analyze and extract the remaining features. Regardless of whether the contract is verified, to facilitate subsequent unified processing, we will perform the decompilation.
Listing 1: Abstract example of a function in contract
FUNCTION name/selector/dest public/private
    CALL name/selector
    JUMP  dest
    DELEGATECALL
    SELFDESTRUCT
@@ 4.2.3. Decompilation
Metadata features Logistic Regression 0.9883 0.6447 0.9245 0.7597 0.0104
Decision Trees 0.9898 0.7500 0.7358 0.7429 0.0050
Random Forest 0.9906 0.7258 0.8491 0.7826 0.0065
MLP 0.9879 0.6400 0.9057 0.7500 0.0104
LightGBM 0.9910 0.7377 0.8491 0.7895 0.0062
Functionality features Logistic Regression 0.9499 0.2778 0.9434 0.4292 0.0500
Decision Trees 0.9883 0.7037 0.7170 0.7103 0.0062
Random Forest 0.9921 0.7963 0.8113 0.8037 0.0042
MLP 0.9906 0.7414 0.8113 0.7748 0.0058
LightGBM 0.9943 0.8654 0.8491 0.8571 0.0027
All features Logistic Regression 0.9894 0.6667 0.9434 0.7812 0.0096
Decision Trees 0.9959 0.9038 0.8868 0.8952 0.0019
Random Forest 0.9970 0.9091 0.9434 0.9259 0.0019
MLP 0.9936 0.7903 0.9245 0.8522 0.0050
LightGBM 0.9981 0.9286 0.9811 0.9541 0.0015
@@
Table 6. The performance metrics for different features and models, evaluated on the complete contract samples.
Specifically, we first decompile the runtime code (deployed on the chain) of the contract to obtain the comprehensive pseudocode. Then, we employ regular expressions to extract information about implemented functions, and delve deeper into parsing the function calls within each identified function. Ultimately, the abstract representation of each function is meticulously documented, exemplified by Listing 1. Regarding public functions, Dedaub’s decompilation process strives to automatically resolve function signatures. As a result, it may either reconstruct the function names or retain the original function selectors, which are unique identifiers derived from the hash of function signatures, used by the EVM to identify and invoke functions. Conversely, for private functions, due to their invocation using jump instructions, they are represented by the destination location where the function resides.
Following this processing, we can acquire various fundamental features in both Implemented Function and Function Call categories, such as PublicFuncNumber, PrivateFuncNumber, ExternalCallNumber, InternalCallNumber, and DelegateCallNumber. Additionally, we can also determine whether the contract can self-destruct.
To glean insights into token-related calls within each function and the implementation of flashloan callbacks, we need to attempt matching each function within the contract with the functions listed in Table 4. It’s worth noting that the function resolved results by decompilation may not always be exhaustive. Therefore, we use both function names and selectors to detect token-related calls and flashloan callbacks. To further extract the MaxTokenCall and AvgTokenCall features, we analyze the execution flow of each public function, which means iteratively traversing the calls within the function, and then counting token-related calls corresponding to the involved private functions and flashloan callbacks.
@@
Furthermore, to ascertain whether the contract is a token contract, we inspect the implemented functions to check if it fully implements all functions of a specific token ERC standard. Specifically, we implement detection for ERC-20 (ERC20, 28), ERC-721 (ERC721, 30), and ERC-1155 (ERC1155, 26) token contract standards. It’s noteworthy that other token ERC standards, such as ERC-777 (ERC777, 31) and ERC-4626 (ERC4626, 29), are backward-compatible with ERC-20. For identifying proxy contracts, despite various proxy patterns, most proxy contracts, following ERC-1967 (ERC1967, 27), typically saves the logic contract’s address in a specific storage slot. Therefore, if a contract involves a delegatecall and the contract bytecode contains the slot value, we consider it a proxy contract, by which the majority of proxy cases are covered.
4.3. Classifier Building
@@
For the fund source feature, which consists of categorical data with low cardinality, we applied one-hot encoding. Additionally, to facilitate better comparability among features of varying dimensions in numerical terms and to improve the accuracy of the classifier, we standardized the data using Z-score normalization. This process centers each feature’s values around 0 with a standard deviation of 1.
@@
Finally, when selecting an appropriate supervised machine learning algorithm to construct the classifier, we explore various models. These models include classical approaches such as Logistic Regression and Decision Trees. We also try ensemble learning approaches, such as Random Forest based on bagging and the LightGBM based on boosting. In addition, we considered the use of a classic neural network architecture, the Multilayer Perceptron (MLP), as one of the candidate models.
5. Evaluation
In this section, we provide a comprehensive exposition of the evaluation results for the built classifier based on our dataset. We first introduce the metrics and methodologies employed for the evaluation in §5.1, then demonstrate the performance of our classifier in §5.2, elucidates the importance of features in §5.3, and finally in §5.4, we conduct a detailed analysis of the failure instances where the classifier made incorrect identifications during testing.
5.1. Metrics and Methodology
@@
In addition to commonly used binary classification performance metrics such as accuracy, precision, recall, F1-Score, we also calculate the False Positive Rate (FPR) metric, which holds particular significance in our scenario. A high FPR implies that the classifier is more likely to incorrectly identify benign contracts as adversarial, potentially leading to significant manual audit costs.
@@
A distinguishing aspect of our research compared to traditional supervised machine learning problems is the explicit chronological relationship between contract deployment. Different contracts deployed over time may exhibit varying characteristics and vulnerabilities, meaning that adversarial contracts may evolve as well. Additionally, our objective is to train a classifier based on existing DeFi incident data to detect potential future attacks. Consequently, traditional cross-validation techniques, which assume each data is independent, may not be suitable for our evaluation.
In response, we adopt a chronological approach to select adversarial contracts for training and testing. We use the earliest 80% of the contracts for training and reserve the latest 20% for testing. Specifically, the 53 most recent adversarial contracts, deployed between August 18, 2023, and December 21, 2023, are allocated for testing, while the remaining 216 earlier contracts are used for training.
Table 7. Comparison of performance with baseline, evaluated on the benign samples excluding token and proxy contracts.
Model Precision Recall F1-Score FPR
LR 0.7246 0.9434 0.8197 0.0315
DT 0.8491 0.8491 0.8491 0.0132
RF 0.9259 0.9434 0.9346 0.0066
MLP 0.8475 0.9434 0.8929 0.0149
LightGBM 0.9273 0.9623 0.9444 0.0066
Forta’s 0.8778 0.5520 0.6208 -
However, since the benign samples were deployed in a different time frame compared to adversarial samples, and the benign contracts are inferred based on the number of interactions, which does not unequivocally confirm a contract’s benign nature, we have adopted a different strategy for them. To maximize the accuracy of the training labels for benign contract, we choose to select the top 80% of benign contracts with the highest interaction for training. This strategy is under the assumption that contracts with more interactions are more likely to be genuinely benign. As a result, 10,400 of the most interacted benign contracts, with the lowest interaction count being approximately 1,300, are used for training. The remaining 2,600 benign contracts with fewer interactions are earmarked for testing.
5.2. Performance
@@
Table 6 presents our test results. Overall, the LightGBM-based classifier outperforms others, demonstrating the best performance across all metrics when using all features. Its F1-Score reaches 0.9541, and it exhibits a remarkably high detection rate for adversarial samples. Out of the 53 adversarial contracts, only one remains undetected, while maintaining an exceptionally low FPR of 0.15%. Other models also perform well. Except the one based on Logistic Regression, the F1-scores of all other classifiers exceed 0.85, indicating the applicability of our approach to the dataset.
Figure 11. Top 15 LightGBM Feature Importance
To the best of our knowledge, this is the first academic work on focused the detection of adversarial contracts utilized in DeFi attack. The only similar research made public so far is by the Forta Network (FortaNetwork, 35), which also uses machine learning. However, their approach differs as it focuses solely on low-level bytecode. We used their method as a baseline for comparison. It is pertinent to mention that Forta excludes token and proxy contracts. Moreover, based on our analysis in  §4.1, it was found that adversarial contracts belonging to the token and proxy types are exceedingly rare. Therefore, to ensure a more reasonable comparison, we evaluate our classifier again by excluding token and proxy contracts from our benign contract samples. The new samples contains a total of 3,020 non-token and non-proxy benign contracts, with the most interacted 2,416 of them used for training and the remainder for testing. The results of the evaluation and comparison are presented in Table 7. It is evident that, while there is a slight decrease in classifier performance compared to using all benign samples, our classifiers still significantly outperform Forta’s most recently published model (released on February 6, 2023) across all performance metrics. This highlights the superiority of our method.
Table 8. Failure instances of the lightGBM-based classifier with a subset of important features. The first contract represents a false negative located on BSC, while the remaining four contracts correspond to false positives on the Ethereum.
Contract Address
Nonce
Fund Source
InputDataLength
GasUsed
Verified
PublicFuncNumber
FlashloanNumber
TokenCallNumber
AvgTokenCall
SelfDestructive
0x0f41f9
3
Anonymous
2616
337467
False
1
0
0
0
False
0xdac3d4
102
Bridge
21060
2195797
False
6
1
10
1.2
True
0x1d2cbc
52
Anonymous
26762
2956727
False
16
0
11
0.75
True
0x70d2f0
0
Anonymous
19336
2099607
False
13
0
12
1.08
True
0xfd594e
1
Anonymous
3306
421261
False
3
0
2
0.67
False
5.3. Feature Importance